### 21 Dec 2023 v1

# Polynomial Time Convergence of the Iterative Evaluation of Datalogo Programs

Abstract
$\text{\sf Datalog}^{\circ}$ is an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative semiring. Like Datalog, $\text{\sf Datalog}^{\circ}$ programs can be evaluated via the natural iterative algorithm until a fixed point is reached. However unlike Datalog, the natural iterative evaluation of some $\text{\sf Datalog}^{\circ}$ programs over some semirings may not converge. It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [10]. Previously, the best known upper bound on the number of iterations until convergence over $p$-stable semirings is $\sum_{i=1}^{n}(p+2)^{i}=\Theta(p^{n})$ steps, where $n$ is (essentially) the output size. We establish that, in fact, the natural iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable semiring converges within a polynomial number of iterations. In particular our upper bound is $O(\sigma pn^{2}(n^{2}\lg\lambda+\lg\sigma))$ where $\sigma$ is the number of elements in the semiring present in either the input databases or the $\text{\sf Datalog}^{\circ}$ program, and $\lambda$ is the maximum number of terms in any product in the $\text{\sf Datalog}^{\circ}$ program.
1 Introduction
Motivated by the need in modern data analytics to express recursive computations with aggregates [10] introduced $\text{\sf Datalog}^{\circ}$, which is an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative semiring. Like Datalog, $\text{\sf Datalog}^{\circ}$ programs can be evaluated via the natural iterative algorithm until a fixed point is reached. The number of iterations until the iterative evaluation of a Datalog program converges is at most the output size. In contrast the natural iterative evaluation of some $\text{\sf Datalog}^{\circ}$ programs over some commutative semirings may not converge. It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [10]. A semiring is $p$-stable if the number of iterations required for any one-variable recursive linear $\text{\sf Datalog}^{\circ}$ program to reach a fixed point is at most $p$, and a semiring is stable if there exists a $p$ for which it is $p$-stable. Previously, the best known upper bound on the number of iterations until convergence is $\sum_{i=1}^{n}(p+2)^{i}=\Theta(p^{n})$ steps, where $n$ is (essentially) the output size, and $p$ is stability of the underlying semiring. In contrast there are no known lower bounds that show that iterative evaluation requires an exponential (in the parameter $n$) number of steps to reach convergence.
Many naturally arising $\text{\sf Datalog}^{\circ}$ programs are linear, which are those where each product only contains at most one output database. [9] showed that iterative evaluation of all linear $\text{\sf Datalog}^{\circ}$ programs converges after $O(pn^{3})$ steps. So iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs over stable commutative semirings converges in a polynomial number of steps. Further, [9] showed that the iterative evaluation of all linear $\text{\sf Datalog}^{\circ}$ programs over commutative $p$-stable commutative semirings containing $L$ elements converges within $O(pn\lg L)$ steps.
The research question we address in this paper is whether the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs can indeed take exponential number of steps to converge. Another way to frame our motivating research question is whether or not polynomial convergence is a special property of linear $\text{\sf Datalog}^{\circ}$ programs that is not shared by general $\text{\sf Datalog}^{\circ}$ programs. Our main finding is stated in Theorem 1.1.
Theorem 1.1.
Let $\bm{S}$ be a $p$-stable commutative semiring. Let $P$ be a $\text{\sf Datalog}^{\circ}$ program where the maximum number of multiplicands in any product is at most $\lambda$. Let $D$ be the input EDB database. Let $\sigma$ be number of the semiring elements referenced in $P$ or $D$. Let $n$ denote the total number of ground atoms in a IDB that at some point in the iterative evaluation of $P$ over $\bm{S}$ on input $D$ have a nonzero associated semiring value. Then the iterative evaluation of $P$ over $\bm{S}$ on input $D$ converges within
$pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)$
steps.
Let us now focus on how we believe one should interpret the statement of Theorem 1.1. It is most natural to focus on how the number of iterations grows as a function of the parameters that are most likely to be large, namely the size of the input databases and the size of the output databases. $\text{\sf Datalog}^{\circ}$ programs, like Datalog programs, use the semiring zero element to signify the nonpresence of a ground tuple in a IDB. Most natural $\text{\sf Datalog}^{\circ}$ programs are monotone in the sense that a ground atom will not be deleted from an IDB after it is added (which means not resetting the associated semiring value to zero after it has become nonzero). For monotone $\text{\sf Datalog}^{\circ}$ programs, $n$ will be the number of ground atoms in the output, and we think it most natural/informative to think of $n$ in this way. Since the stability parameter $p$ is only dependent on the underlying semiring, and is independent of the input database and even in some sense independent of the $\text{\sf Datalog}^{\circ}$ program, we think it is best to view $p$ as a constant that does not scale with the input size or output size. It is natural to think of the program $P$ as being modestly sized, so it’s natural to also think of $\lambda$ as a constant that does not scale with the input size or output size. It’s natural that as the size of the input database scales that there will be significantly more semiring elements referenced in the input database than in the $\text{\sf Datalog}^{\circ}$ program. Thus it is natural to view $\sigma$ as essentially upper bounded by the number of elements in the input databases. Under these assumptions, Theorem 1.1 can be interpreted to say that the number of iterations is
$O(pn^{4}\sigma(\lg\lambda+\lg\sigma))$ (1)
So if one is in a situation where the size of the output databases are quite small compared to the size of the input databases, then the worst-case number of iterations grows at a rate at most (essentially) linearly with the size of the input databases. Otherwise, if one is in the situation where the size of the output databases are quite large compared to the size of the input database (or the input databases have a lower diversity of semiring values), then the worst case number of iterations grows proportionally to the fourth power of the size of the output databases. Finally, if one is in the situation where the size of the output databases are comparable to the size of the input databases (and the input databases have a wide diversity of semiring values), then the worst case number of iterations grows proportionally to the fifth power of the size of the output (or input) databases. Recall that [9] showed that in the worst case the number of iterations for the iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs grows proportional to the cube of the size of the output database.
Let us assume for a moment that the number $L$ of elements in the semiring is large. Then to the extent that one can reasonably interpret $n^{4}$ (or maybe $n^{5}$) as being not too much worse than $n^{3}$, Theorem 1.1 can reasonably be interpreted as saying that the worst-case time to convergence for the iterative evaluation of general $\text{\sf Datalog}^{\circ}$ programs is not too much worse than the time to convergence for the iterative evaluation of linear $\text{\sf Datalog}^{\circ}$ programs. On the other hand, if $L$ is not too large, then it is reasonable to interpret the $O(pn\lg L)$ convergence time for linear $\text{\sf Datalog}^{\circ}$ programs in [9] to be significantly better than the bound for general $\text{\sf Datalog}^{\circ}$ programs in Theorem 1.1.
In the rest of the introduction we give a fast paced coverage of the most essential background knowledge required to understand this result ( For a deeper dive see  [10]). We then give a brief technical overview of the proof of Theorem 1.1.
1.1 Background
1.1.1 Datalog
A (traditional) Datalog program $P$ consists of a set of rules of the form:
[…]
where $R_{0},\ldots,R_{m}$ are predicate names (not necessarily distinct) and each $\bm{X}_{i}$ is a tuple of variables and/or constants. The atom $R_{0}(\bm{X}_{0})$ is called the head, and the conjunction $R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{m}(\bm{X}_{m})$ is called the body. Multiple rules with the same head are interpreted as a disjunction. A predicate that occurs in the head of some rule in $P$ is called an intensional database predicate (IDB), otherwise it is called an extensional database predicate (EDB). The EDBs form the input (databases), and the IDBs represent the output (databases) computed by the Datalog program. The finite set of all constants occurring in an EDB is called the active domain, and denoted ADom. There is an implicit existential quantifier over the body for all variables that appear in the body, but not in the head, where the domain of the existential quantifier is ADom. Thus, a Datalog program can also be viewed as a collection of unions of conjunctive queries (UCQs), one UCQ for each IDB.
Example 1.
One classic example of a Datalog program is the following one, which computes the transitive closure of a directed graph, defined by the edge relation $E(X,Y)$:
[…]
Here $E$ is an EDB predicate, $T$ is an IDB predicate, and ADom consists of the vertices in the graph. Written as a UCQ, where the quantifications are explicit, this program is:
[…]
A Datalog program can be thought of as a function, called the immediate consequence operator (ICO), where the ICO adds a ground atom $A$ to a IDB $R$ if there a predicate where $R$ is the head, and $R(A)$ is logically implied by the body. The iterative evaluation of a Datalog program works in rounds/steps, where on each round the ICO is applied to the current state. In the context of the Datalog program in Example 1, if $(a,b)$ was a ground atom in $T$, meaning that it known that there is a directed path from vertex $a$ to vertex $b$, and $(b,c)$ was a ground atom in $E$, meaning that there is a directed edge from $b$ to $c$, then the ICO would add the ground atom $(a,c)$ to $T$ (if it was not already present). This is because the Datalog program states that $T(a,c)$ is logically implied by $T(a,b)$ and $E(b,c)$.
1.1.2 $\text{\sf Datalog}^{\circ}$
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the union of conjunctive queries are now replaced with sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$. Each rule has the form:
[…]
where sum is over the active ADom of the variables not in $X_{0}$. Further each ground atom in an EDB predicate or IDB predicate is associated with an element of the semiring $\bm{S}$, and the element associated with a tuple in an EDB predicate is specified in the input. The EDBs form the input database, and the ground atoms in the IDB’s that have an associated semiring value that is nonzero represent the output instance computed by the $\text{\sf Datalog}^{\circ}$ program. Note that in a Datalog program the ground atom present in one of the databases can be thought of as those that are associated with the element $1$ in the standard Boolean semiring.
Example 2.
One important class of a $\text{\sf Datalog}^{\circ}$ programs is those of the following form:
[…]
Here $E$ is an EDB predicate, and $T$ is an IDB predicate. Note that this $\text{\sf Datalog}^{\circ}$ program is the Datalog program given in line (3) with $\vee$ replaced by $\oplus$, $\wedge$ replaced by $\otimes$ and $\exists_{Z}$ replaced by $\bigoplus_{Z}$. Here $\oplus$ is the semiring addition operation, and $\otimes$ is the semiring multiplication operation and $\bigoplus_{Z}$ is aggregation, that is an iterative application of $\oplus$ over ADom. Note that when interpreted over the Boolean semiring, this $\text{\sf Datalog}^{\circ}$ program is essentially the transitive closure program from Example 1.
Example 3.
When interpreted over the tropical semiring $\text{\sf Trop}^{+}=(\mathbb{R}_{+}\cup\{\infty\},\min,+,\infty,0)$, the $\text{\sf Datalog}^{\circ}$ program in Example 2 solves the classic All-Pairs-Shortest-Path problem (APSP), which computes the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$.
[…]
A $\text{\sf Datalog}^{\circ}$ program can be thought of as an immediate consequence operator (ICO). To understand how the ICO works in $\text{\sf Datalog}^{\circ}$, consider a rule with with head $R$ and let $A$ be a ground atom for $R$ with associated semiring value $y$, and assume that for $A$ the body of this rule evaluates to the semiring element $x$. As a result of this, the ICO adds $x$ to the semiring value associated with $A$, resulting in a new value of $x+y$. Note that the functioning of the $\text{\sf Datalog}^{\circ}$ ICO, when the semiring is the standard Boolean semiring, is identical to how the Datalog ICO functions. The iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program works in rounds/steps, where initially the semiring element 0 is associated with each possible ground atom in an IDB, and on each round the ICO is applied to the current state. So in the context of the $\text{\sf Datalog}^{\circ}$ program in Example 3, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $E$ with associated semiring element $y$, meaning that there is a directed edge from $b$ to $c$ with length $y$, then the ICO would make the semiring element associated with $A$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
A $\text{\sf Datalog}^{\circ}$ program is linear if every rule has no more than one IDB prediciate in its body. The $\text{\sf Datalog}^{\circ}$ program in Example 2 is linear. While many natural $\text{\sf Datalog}^{\circ}$ programs are linear, there are also natural nonlinear $\text{\sf Datalog}^{\circ}$ programs.
Example 4.
As a classic example of a nonlinear $\text{\sf Datalog}^{\circ}$ program, consider the following alternate program over the tropical semiring for computing the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$:
[…]
In the $\text{\sf Datalog}^{\circ}$ program in Example 4, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $T$ with associated semiring element $y$, meaning that the shortest known directed path from vertex $b$ to vertex $c$ has length $y$, then the ICO would make the semiring element associated with $(a,c)$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
1.1.3 Stability
Given a semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ and $u\in S$, let $u^{(p)}:=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}$, where $u^{i}:=u\otimes u\otimes\cdots\otimes u$ ($i$ times). Then $u\in S$ is $p$-stable if $u^{(p)}=u^{(p+1)}$, and semiring $\bm{S}$ is $p$-stable if every element $u\in S$ is $p$-stable. A function $f:S^{n}\to S^{n}$ is $p$-stable if $f^{(p+1)}(\bm{0})=f^{(p)}(\bm{0})$, where $\bm{0}$ is the all zero vector, and $f^{(k)}$ is the $k$-fold composition of $f$ with itself. The stability index of $f$ is the smallest $p$ such that $f$ is $p$-stable. The convergence time of a $\text{\sf Datalog}^{\circ}$ program is the stability index of its ICO. It is important to note that a semiring is $p$-stable if and only if every linear function $f:S\to S$ is $p$-stable.
1.1.4 Grounding the ICO
Since the final associated semiring values of the ground atoms in an IDB are not initially known, it is natural to think of them as (IDB) variables. Then the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$program is a map $\bm{f}:S^{n}\rightarrow S^{n}$, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program. For instance, in (6), there would be one variable for each pair $(x,y)$ of vertices where there is a directed path from $x$ to $y$ in the graph. So the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program has the following form:
$-\displaystyle X_{1}\text{ :- }$- $f_{1}(X_{1},\ldots,X_{n})$
$\ldots$ (8)
$-\displaystyle X_{n}\text{ :- }$- $f_{n}(X_{1},\ldots,X_{n})$
where the $X_{i}$’s are the IDB variables, and $f_{i}$ is the component of $\bm{f}$ corresponding to the IDB variable $X_{i}$. Note that each component function $f_{i}$ is a multivariate polynomial in the IBD variables of degree at most the maximum number of terms in any product in the body of some rule in the $\text{\sf Datalog}^{\circ}$ program. After $q$ iterations of the iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program, the semiring value associated with the ground atom corresponding to $X_{i}$ will be:
${f}_{i}^{(q)}(\bm{0})$ (9)
1.1.5 Context Free Languages
It is convenient to reason about the formal expansion of $f_{i}^{(q)}(\bm{0})$ using context-free languages (CFL). To explain this, it is probably best to start with a concrete example.
Example 5.
The following map $\bm{f}=(f_{1},f_{2})$:
$\begin{bmatrix}A\\ B\end{bmatrix}\to\begin{bmatrix}aAB+bB+c\\ cAB+bA+a\end{bmatrix}$ (14)
can be represented by the following context free grammar $G$:
$A\to aAB\ |\ bB\ |\ c$ $B\to cAB\ |\ bA\ |\ a$
More generally the variables in $f$ become non-terminals in $G$, the constants in $f$ become the terminals in $G$, multiplication in $f$ becomes concatenation in $G$, and addition in $f$ becomes the or operator $\mid$ in $G$. Given a parse tree $T$ for the grammar, define the yield $Y(T)$ of $T$ to be the string of terminal symbols at the leaves of $T$, and the product yield $Z(T)$ to be the product of the semiring values in $Y(T)$. Let $\mathcal{T}^{i}_{q}$ denote the set of all parse trees with starting non-terminal $X_{i}$, and depth $\leq q$. Note that then:
$f_{i}^{(q)}(0)$ $=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T).$ (15)
That is, the value of the semiring value associated with a ground atom after in a IDB after $q$ iterations is the sum over the product of the leaves of parse trees of depth at most $q$ and rooted at $X_{i}$.
1.1.6 Parikh’s Theorem
The upper bound on the time to convergence for iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs in  [10] essentially relied on black-box application of Parikh’s theorem [19]. The Parikh image of a word $w\in\Sigma^{*}$, denoted by $\Psi(w)$, is the vector $\Psi(w)=(k_{1},\dots,k_{\sigma})\in\mathbb{N}^{\sigma}$ where $k_{i}$ is the number of occurrences of the letter $a_{i}\in\Sigma$ that occur in the word $w$ (So $|\Sigma|=\sigma$). Then using our assumption that the underlying semiring $\bm{S}$ is commutative, we observe that
$f_{i}^{(q)}(\bm{0})$ $=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T)=\sum_{T\in\mathcal{T}^{i}_{q}}\prod_{j=1}^{\sigma}a_{j}^{\Psi_{j}(Y(T))}$ (16)
where $\Psi_{j}(w)$ is component $j$ of the Parikh image. One version of Parikh’s theorem [19] states that the Parikh images of the words in in a context free language forms a semi-linear set. A set is then semi-linear if it is a finite union of linear sets. A set $\mathcal{L}\subseteq\mathbb{N}^{\sigma}$ is said to be linear if there exist offset vector $\bm{v}_{0}$ and basis vectors $\bm{v}_{1},\ldots,\bm{v}_{\ell}\in\mathbb{N}^{\sigma}$ such that $\mathcal{L}$ is the span of of these vectors, that is if:
$\mathcal{L}=$$\{{\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}}\mid{k_{1},\ldots,k_{\ell}\in\mathbb{N}}\}.$
If a vector
$\bm{v}=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$
then we say $(k_{1},\ldots,k_{\ell})$ is a linear representation of $v$ within $\mathcal{L}$.
The textbook proof of Parikh’s theorem (see [12]) uses what we will call a wedge. A wedge within a parse tree $T$ can be specified by identifying two internal nodes in the parse tree that correspond to the same nonterminal, say $A$, and that have an ancestor-descendent relation. The corresponding wedge $W$ then consists of the nodes in the parse tree that are descendents of the top $A$, but not the bottom $A$. See Figure 1.
Figure 1: Illustration of a wedge $W$
1.2 Technical Overview
We now give a technical overview of the proof of Theorem 1.1. This proof has three logical parts. The first part, which is stated in Theorem 1.2, strengthens some aspects of the standard statement of Parikh’s theorem [12].
Theorem 1.2.
Let $L$ be a context free language generated by a grammar $G=(N,\Sigma,R,S)$, where $N$ is the collection of nonterminals, $\Sigma$ is the collection of terminals, $R$ is the collection of rules, and $S\in N$ is the start non-terminal. Let $n$ be the cardinality of $N$ and let $\lambda$ be the maximum number of symbols on the righthand side of any rule. Then there exists a finite semi-linear set $\mathcal{M}$ with the following properties:
  1. 1.
    A word $w$ is in $L$ if and only if $\Psi(w)$ is in some linear set $\mathcal{L}\in\mathcal{M}$.
  2. 2.
    Every linear set $\mathcal{L}\in\mathcal{M}$ has an associated offset vector $v_{0}$ and basis vectors $v_{1},\ldots v_{\ell}$, with the property that :
    1. (a)
      There is a word $w\in L$ such $\Psi(w)=v_{0}$ and $w$ can be generated by a parse tree with depth at most $n(n+3)/2$.
    2. (b)
      Each basis vector has an associated wedge with depth at most $n(n+3)/2$.
    3. (c)
      The 1-norm of the offset vector, and each basis vector, is at most $\lambda^{n(n+3)/2}$.
    4. (d)
      For each vector $\bm{v}=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$ in the span of $\mathcal{L}$ it is the case that there is a word $w\in L$, with $\Psi(w)=\bm{v}$, where $w$ can be generated by a parse tree with depth at most $(k+1)n(n+3)/2$, where $k=k_{1}+k_{2}+\ldots+k_{\ell}$.
The proof of Theorem 1.2 is given in Section 2. The most important way that Theorem 1.2 extends the standard version of Parikh’s theorem is property 2(d), which upper bounds the depth of some parse tree of a word by the 1-norm of the representation of that word. The bound given in the textbook proof [12] of Parikh’s theorem gives a depth bound that is exponentially large (which essentially yields the exponential upper bound on the convergence time [10]). To achieve property 2(d), our proof contains a constructive forward process $\mathcal{P}$ that creates a linear set $\mathcal{L}\in\mathcal{M}$ from the parse tree $T$ of some word $w\in L$ by removing wedges from $T$. We are careful to design $\mathcal{P}$ so that it is reversible, that is to recover a parse tree from a linear representation we can just reverse the process $\mathcal{P}$. To accomplish this we need that in the forward process every wedge that is removed does not remove any nonterminal from the parse tree. Our process $\mathcal{P}$ insures that the depth of the parse tree for the offset and the wedges for the basis vectors have height $O(n^{2})$.
The second part of the proof of Theorem 1.1 is stated in Lemma 1.3, which states that for every word $w\in L$ there exists a linear set $\mathcal{L}$ in $\mathcal{M}$ such that $w\in\mathcal{L}$ and has a linear representation with respect to the basis vectors of $\mathcal{L}$ that has small support.
Lemma 1.3.
Let $L$ be an arbitrary CFL, and $\mathcal{M}$ the semilinear set that is guaranteed to exist in Theorem 1.2. Then for every word $w\in L$ there exists a linear set $\mathcal{L}$ in $\mathcal{M}$ such that $w\in\mathcal{L}$ and $w$ has a linear representation with respect to the basis vectors in $\mathcal{L}$ that has support size at most:
$2(\sigma(n(n+3)/2)(\lg(\lambda+1)+4\sigma\lg\sigma)$
The proof of Lemma 1.3, given in Section 3, uses properties of $\mathcal{L}$ established in our strengthened version of Parikh’s theorem (Theorem 1.2), and the pigeon hole principle to establish that any word in $\mathcal{L}$ that has a linear representation with respect to $\mathcal{L}$ with large support, also has a linear representation with smaller support. In particular we use the finding that all the offset and basis vectors have (relatively) small 1-norms. This makes formal the intuition that one might draw from standard vector spaces that the number of basis vectors needed to represent a vector/word in the span of some basis vectors shouldn’t be more than the dimensionality of the spanned space.
Finally in Section 4 we use Lemma 1.3 to prove Theorem 1.1. The proof considers the context free language $L$ corresponding to a component $i$ of $f$, and then an arbitrary word $w\in L$. By Lemma 1.3 it must be the case that the Parikh image of $w$ has a representation
$\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$
in some linear set $\mathcal{L}$ of the semilinear set $\mathcal{M}$ that has small support. If each coefficient $k_{i}$ is small, then by condition 2(d) in Theorem 1.2 it must be the case that there is a low depth parse tree $T$ for $w$, and thus the summand corresponding to $w$ will be added to semiring value associated with $X_{i}$ in a relatively small number of steps in the naive evaluation of the $\text{\sf Datalog}^{\circ}$ program.  Otherwise assume some coordinate, say without loss of generality that $k_{\ell}$ is large. Then note that each of the Parikh images
$\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+(k_{\ell}-j)\bm{v}_{\ell}$
for $j\in[1,p]$ must have been added to the semiring value associated with $X_{i}$ before the summand associated with $w$ was added, and thus by the stability of the product corresponding to $v_{\ell}$, when the Parikh image of $w$ is added to the semiring value associated with $X_{i}$, and thus it must be the case that this semiring value will not change.
2 The Strengthened Parikh’s Theorem
Our goal in this section is to prove Theorem 1.2. A word’s depth is defined as the minimum depth of a parse tree that yields the word. Similarly, we say that a vector $v$’s depth is defined as the minimum depth of a word $w$ with $\Psi(w)=v$.
We begin our proof by defining a semi-linear set $\mathcal{M}$ with the desired properties. Let $c:=n(n+3)/2$ throughout this section. Let $\mathcal{T}^{S}_{c}$ denote the set of all parse trees (starting with the non-terminal $S$) of depth at most $c$. Recall that $T$’s yield, denoted as $Y(T)$, is the word obtained by a parse tree $T$. For a wedge $W$, $Y(W)$ is analogously defined by ignoring the unique non-terminal leaf node in the wedge $W$. Let $N(T)$ denote the set of non-terminals appearing somewhere in $T$. Let $\mathcal{W}^{A}_{c}(T)$ be the collection of wedges that have a non-terminal $A$ as the root, have non-terminals only from $N(T)$ and have depth (or equivalently height) at most $c$. Let $\mathcal{B}_{c}(A,T):=\{Y(W)\;|\;W\in\mathcal{W}^{A}_{c}(T)\}$. For notational brevity, we may use $\mathcal{B}(A,T)$ instead of $\mathcal{B}_{c}(A,T)$.
Then, for each tree $T$ in $\mathcal{T}^{S}_{c}$, we define a linear set where the offset vector is $\Psi(Y(T))$ and the basis vectors are $\cup_{V\in N(T)}\Psi(\mathcal{B}(V,T))$. Here, $\Psi(L^{\prime})$ denotes the collection of vectors corresponding to the subset of words, $L^{\prime}$. Notice that because of the way we created the offset vector and basis vectors, there is a parse tree in $\mathcal{T}^{S}_{c}$ corresponding to the offset vector and a wedge corresponding to each basis vector, all of depth at most $c$.
In the following we recall the definition of wedges (Figure 1) and define how to index them. For an arbitrary parse tree $T$ we will map it to a tree in $\mathcal{T}^{S}_{c}$ by iteratively removing a wedge.
Definition 2.1.
Define a wedge of a parse tree $T$ as follows. Consider two occurrences of a non-terminal $A$ in $T$ where one is an ancestor of the other. Let $A^{\prime}$ be the ancestor node and $A^{\prime\prime}$ the descendant node. The wedge induced by the pair $(A^{\prime},A^{\prime\prime})$ is defined as the subtree rooted at $A^{\prime}$ with the subtree rooted at $A^{\prime\prime}$ removed. The wedge is denoted as $W(A^{\prime},A^{\prime\prime})$. The wedge’s depth is defined as the maximum number of edges from $A^{\prime}$ to a leaf node in $W(A^{\prime},A^{\prime\prime})$.
We would like to keep the following invariant, throughout the iterative process.
Figure 2: The right tree $T$ is recovered from the left tree $T^{\prime}$ by augmenting the wedge $W$.
Lemma 2.2.
Given a parse tree $T$ of depth greater than $c$ starting with $S$, we can obtain a parse tree $T^{\prime}$ starting with $S$ that satisfies the following:
  1. 1.
    (Preserving Non-terminals) $N(T)=N(T^{\prime})$.
  2. 2.
    (Reversibility) $T$ can be obtained by replacing one non-terminal leaf $A$ in $T^{\prime}$ with a wedge $W\in\mathcal{W}_{c}^{A}(T^{\prime})$ for some $A\in N(T^{\prime})$.
Alternatively, the second property means that $T$ can be obtained from $T^{\prime}$ by augmenting $T^{\prime}$ with a wedge $W$ of depth at most $c$ corresponding to a vector in $\Psi(\mathcal{B}(A,T^{\prime}))$ for some non-terminal $A$ in $N(T^{\prime})$. Here it is worth noting that $\mathcal{W}^{A}_{c}(T)=\mathcal{W}^{A}_{c}(T^{\prime})$ because $N(T)=N(T^{\prime})$. See Figure 2 for an illustration of reversibility.
The first property in the lemma is worth special attention. Suppose we obtained $T^{\prime}$ from $T$ by repeatedly applying the lemma, but without guaranteeing the first property. Suppose $\{W_{i}\}$ are the wedges we removed in the process; so $W_{i}$ is the wedge removed in iteration $i$. Then, we may not be able to augment $T^{\prime}$ with an arbitrary subset of the wedges, which is critical to establish the “if” direction of the first property of Theorem 1.2.
To show Lemma 2.2, consider an arbitrary parse tree $T$ of depth more than $c$. We show how to obtain $T^{\prime}$ by collapsing a wedge induced by two occurrences of the same non-terminal. Below, we describe how we find a “good” pair of two occurrences of the same non-terminal we want to collapse. We first define what makes pairs good in the following.
Definition 2.3.
For a given parse tree $T$, we say a pair of ascendant and descendant nodes $(A^{\prime},A^{\prime\prime})$ of the same non-terminal $A$ is good if it satisfies the following:
  • Let $T^{\prime}$ be the tree $T$ with the wedge $W(A^{\prime},A^{\prime\prime})$ removed. We have $N(T)=N(T^{\prime})$.
  • The height of $A^{\prime}$ is at most $c$. In other words, the subtree rooted as $A^{\prime}$ has depth at most $c$.
In the following we will show that a tree of large depth must have a good pair. Note that we will immediately have Lemma 2.2 as corollary if we prove the following the lemma.
Lemma 2.4.
A parse tree $T$ of depth at least $c$ has a good pair of nodes.
Proof.
We prove the lemma by an induction on the number of non-terminals. Consider an arbitrary node $v$ of the largest depth, which must be at least $c=(n+3)n/2=(n+1)+n+\ldots+2$. Since the base case $n=1$ is trivial, suppose $n\geq 2$. Consider the unique path from $v$ to the root $S$ in $T$. The height of a node $u$ on the path is defined as the number of nodes below $u$ on the path, including $u$.
By walking up the path, we find the first two occurrences of the same non-terminal $V_{1}$. Say they appear at nodes $v(V_{1})$ and $u(V_{1})$, where $u(V_{1})$ is an ascendant of $v(V_{1})$. Note that $u(V_{1})$ has height at most $n+1$ due to the pigeon hole’s principle. Thus, the two nodes have heights differing by at most $n$.
If $(u(V_{1}),v(V_{1}))$ is a good pair, we are done. If not, it means that the wedge $W(u(V_{1}),v(V_{1}))$ must include a non-terminal that doesn’t appear anywhere else in the tree. Consequently, we now consider the tree $T_{2}$ with the subtree rooted at $u(V_{1})$ removed. Note that the subtree $T_{2}$ has at most $n-1$ non-terminals and has depth at least $n+(n-1)+\ldots+2$. By induction, this implies that $T_{2}$ must have a good pair. It is easy to see that the pair remains to be good with respect to $T$ as well. ∎
We are now ready to prove Theorem 1.2.
Property (1) “only if” direction. Given a parse tree $T$ for $w\in L$, suppose we obtained a sequence of trees $T_{0}=T,T_{1},\ldots,T_{\eta}$ by repeatedly applying Lemma 2.2, where $T_{\eta}\in\mathcal{T}^{S}_{c}$ and $T_{i}$ is obtained from $T_{i-1}$ by deleting a wedge $W_{i}$ in $\mathcal{W}^{A}_{c}(T)$ for some non-terminal $A$ in $N(T)=N(T_{\eta})$; note $\mathcal{W}^{A}_{c}(T)=\mathcal{W}^{A}_{c}(T_{\eta})$ since $N(T)=N(T_{1})=\ldots=N(T_{\eta})$. Let $b_{i}=\Psi(Y(W_{i}))$. Clearly, $\Psi(w)$ can be expressed as $\Psi(Y(T_{\eta}))+\sum_{i=1}^{\eta}b_{i}$.
Since we created a linear set for each parse tree in $\mathcal{T}^{S}_{c}$, thus for $T_{\eta}$, this is a linear representation within $\mathcal{L}$ which consists of offset vector $\Psi(Y(T_{\eta}))$ and basis vectors $\cup_{V\in N(T_{\eta})}\Psi(\mathcal{B}(V,T_{\eta}))$. This proves the “only if” part of the first property.
Property (1) “if” direction and Property (2)(d). Conversely, suppose $w$ has a linear representation within some $\mathcal{L}\in\mathcal{M}$. Say the linear representation is $\Psi(Y(T^{\prime}))+\sum_{i=1}^{k}b_{i}$ for some $T^{\prime}\in\mathcal{T}_{short}$. Note that for each basis vector $b_{i}$, there exists $V_{i}\in N(T^{\prime})$ such that $b_{i}\in\Psi(\mathcal{B}(V_{i},T^{\prime}))$. Because of the way we defined linear sets, $b_{i}=\Psi(Y(W_{i}))$ for some $W_{i}\in\mathcal{W}_{c}^{V_{i}}(T^{\prime})$. We can augment $T^{\prime}$ with the wedges $W_{i}$ in an arbitrary order. Adding each wedge $W_{i}$ increases the tree depth by at most $c$. By repeating this for each $b_{i}$, we obtain a parse tree corresponding to the given linear representation and its depth is at most $c(k+1)$.
Property (2)(a,b,c). By definition of the offset and basis vectors, it immediately follows that their depth is at most $c$. Furthermore, the 1-norm of any of them is at most $\lambda^{c}$ because each node has at most $\lambda$ children.
3 Small Support Representations
In this section we prove Lemma 1.3. Recall that this lemma states that the Parikh image of every word in a context-free language $L$ has a linear representation with small support.
Proof.
To streamline our analysis, we will assume that $\lambda\geq 2$. Otherwise, we can pretend that this is the case by adding an unused terminal to $\Sigma$. Then, we will only need to increment the value of $\lambda$ by one in the final bound.
Consider an arbitrary word $w$ in our language $L$. Let $\mathcal{L}$ be a linear set that contains $w$. Since $\mathcal{L}$ is linear it can be defined as follows:
$\mathcal{L}=$$\{{\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}}\mid{k_{1},\ldots,k_{\ell}\in\mathbb{N}}\}$
For any word $w$, by renumbering the basis vectors we can assume that basis vectors $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{h}$ have non-zero coefficients $k_{1},k_{2},\ldots,k_{h}>0$. So we have:
$w=$ $\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{h}\bm{v}_{h}$
[…]
$\sum_{i\in H_{1}}\bm{v}_{i}=\sum_{i\in H_{2}}\bm{v}_{i}$ (17)
We will prove this claim by upper-bounding
$K:=|\{\sum_{i\in H}\bm{v}_{i}\;|\;H\subseteq[h]\}|,$
the number of distinct vectors we can generate by choosing a subset of vectors from $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{h}$ (we choose each basis vector up to one copy).
Let $M$ be the maximum 1-norm of any $\bm{v}_{i}$, i.e., $||\bm{v}_{i}||_{1}\leq M$, for $i\in[h]$. Thus, we have $||\sum_{i\in H}\bm{v}_{i}||_{1}\leq Mh$ for any $H\subseteq[h]$. We use the following well-known fact: the number of distinct vectors in $Z^{d}_{\geq 0}$ with 1-norm of $k$ is exactly (${k+d-1\choose d-1}\leq(k+d-1)^{d-1}$. In our case, $k\leq Mh$ and $d=\sigma$. Thus, we have $K\leq(Mh+\sigma-1)^{\sigma-1}(Mh+1)\leq(Mh+\sigma)^{\sigma}$. If $\sigma=1$ (there exits only one terminal), we have a tighter bound of $K\leq M+(M-1)+\ldots+M-(h-1)=h(2M-(h-1))/2$.
We aim to show
$2^{h}>M,$ (18)
which will imply the existence of a desired pair of $H_{1}$ and $H_{2}$ satisfying Eqn. (17) due to the pigeon hole principle.
We first consider the case when $\sigma\geq 2$. We shall now establish
$2^{h}>(Mh+\sigma)^{\sigma}\textnormal{ when $\sigma\geq 2$}$ (19)
By taking the logarithm, the inequality (19) is equivalent to:
$h>\sigma\lg(Mh+\sigma)$ (20)
We know from our main Theorem 1.2 that $M\leq\lambda^{n(n+3)/2}$. Thus we know that the following equation would imply Eqn. (20):
$h>\sigma\lg(h\lambda^{n(n+3)/2}+\sigma)$ (21)
Since $\lg x$ is sub-additive when $x\geq 2$, using the assumptions that $\lambda,\sigma\geq 2$, it is sufficient to show:
$\;\;h>\sigma\lg(h\lambda^{n(n+3)/2})+\sigma\lg\sigma=\sigma\lg h+\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
$\Leftrightarrow$ $\;\;h-\sigma\lg h>\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
We show that
$h-\sigma\lg h\geq h/2$
when $h\geq 8\sigma\lg\sigma$.
Since $h/2-\sigma\lg h$ is increasing in $h$ when $h\geq 8\sigma\lg\sigma$, we have $h/2-\sigma\lg h\geq 4\sigma\lg\sigma-\sigma\lg(8\sigma\lg\sigma)=\sigma\lg(\sigma^{3}/(8\lg\sigma))\geq 0$ when $\sigma\geq 2$, as desired. Thus, we have shown that $2^{h}>K$ when $\sigma\geq 2$.
If $\sigma=1$, as mentioned before, we have $K\leq M+(M-1)+\ldots+M-(h-1)=h(2M-(h-1))/2$. Thus, we have:
$2^{h}>K$
[…]
which is true due to the assumption that $h>2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma)=n(n+3)\lg\lambda$; recall $\sigma=1$. Thus, for all values of $\sigma$, we have shown that $2^{h}>K$, which implies the existence of $H_{1}\neq H_{2}\subseteq[h]$ satisfying Eqn. (17).
Now that we have established the existence of a desired pair of $H_{1}$ and $H_{2}$, we now explain how to construct a new representation of $w$ that contains less basis vectors. First observe that one of two sets $H_{1},H_{2}$ doesn’t contain the other since no basis vectors are $\bm{0}$ and we have Eqn. (17). Let $i$ be $\arg\min_{i^{\prime}\in H_{1}-H_{2}}k_{i^{\prime}}$, breaking ties arbitrarily. Then for $j\in H_{1}-H_{2}$ let $k_{j}^{\prime}=k_{j}-k_{i}$, for $j\in H_{2}-H_{1}$ let $k_{j}^{\prime}=k_{j}+k_{i}$, and for all other $i$ let $k_{j}^{\prime}=k_{j}$.
Note that there was no change in the sum, i.e.,
$w=$ $\;\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{h}\bm{v}_{h}$
$=$ $\;\bm{v}_{0}+k^{\prime}_{1}\bm{v}_{1}+\ldots+k^{\prime}_{h}\bm{v}_{h}$
However, this new representation $\langle k^{\prime}_{1},k^{\prime}_{2},\ldots,k^{\prime}_{h}\rangle$ has a strictly smaller support since $k^{\prime}_{i}=0$. This contradicts the minimality in the definition of $h$.
Thus, we have shown that $w$ has a linear representation of support size at most $2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma)$ assuming $\lambda\geq 2$. To remove this assumption we add an unused terminal to $\Sigma$, which increments the value of $\lambda$ by one in the bound. ∎
4 Bounding the Number of Iterations
This section is devoted to proving Theorem 1.1.
Consider an arbitrary IDB variable $X_{i}$ and let $L$ be the CFL associated with this variable. For the sake of analysis we impose a linear order on the words $w\in L$. Let $\mathcal{M}$ be a semi-linear set that satisfies the properties stated in Theorem 1.2. Let $\ell(w)$ be defined as the minimum 1-norm of all possible linear representations of $w$ in terms of the offset and basis vectors of some $\mathcal{L}\in\mathcal{M}$. We order the words $w$ by increasing order of $\ell(w)$ breaking ties arbitrarily. We will show adding a summand $\Psi(w)$ with $\ell(w)>ph$ doesn’t change the following summation,
$\sum_{\ell=1}^{\ell(w)-1}\sum_{w^{\prime}:\ell(w^{\prime})=\ell}\Psi(w^{\prime}),$ (22)
where
$h=2(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma)$
Consider an arbitrary $w\in L$ with $\ell(w)>ph$. We know from Lemma 1.3 that there is a sparse linear representation
$\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{\gamma(1)}+\ldots+k_{h}\bm{v}_{\gamma(h)}$
of the Parikh image of $w$ in terms of the offset and basis vectors for some linear set $\mathcal{L}\in\mathcal{M}$. Since $\ell(w)>ph$, by the pigeon hole principle there must exist an $j\in[h]$ such that $k_{j}\geq p+1$. We can assume without loss of generality (by renumbering the basis vectors) that $j=1$. Now consider the Parikh images:
$\bm{v}_{0}+(k_{1}-g)\bm{v}_{\gamma(1)}+\ldots+k_{h}\bm{v}_{\gamma(h)}$
for $g\in[1,p+1]$. By the first property of Theorem 1.2 there are $p$ words $w_{1},\ldots,w_{p}$ in $L$ that have these Parikh images. Further, since for all $i\in[1,p+1]$ $\ell(w_{i})<\ell(w)$ values, each $\Psi(w_{i})$ appears in Eqn. (22). But then by the $p$-stability of the element corresponding to $\bm{v}_{\gamma(1)}$, adding the summand $\Psi(w)$ doesn’t change the sum.
Then by condition 2(d) of Theorem 1.2, we know that any word $w\in L$ with $\ell(w)\leq ph$ can be yielded by a parse tree of depth at most
$(ph+1)n(n+3)/2$
which substituting in the value of $h$ is upper bounded by
$q:=pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)$ (23)
Thus, we have shown that :
$f_{i}^{(q)}(\bm{0})=f_{i}^{(q+1)}(\bm{0})$
where
$f_{i}^{(q)}(\bm{0})=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T).$
and $\mathcal{T}^{i}_{q}$ denotes the collection of the parse trees of depth at most $q$ starting with $X_{i}$. This completes the proof.
5 Related Work
If the semiring is naturally ordered111 $\bm{S}$ is naturally ordered if the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$ is a partial order. , then the least fixed point of a $\text{\sf Datalog}^{\circ}$ program is the least fixed point of $f$ under the same partial order extended to $S^{n}$ componentwise. This is the least fixed point semantics of a $\text{\sf Datalog}^{\circ}$ program. Under (ground-atom) set containment, this function has a unique least fixed point which can be obtained via repeatedly applying the natural iterative evaluation algorithm that applies ICO until a fixed point is reached [1]. This least fixed point is the semantics of the given Datalog program.
Computing the least fixed point solution to a recursive $\text{\sf Datalog}^{\circ}$ program boils down to solving fixed point equations over semirings. In particular, we are given a multi-valued polynomial function $f:S^{n}\to S^{n}$ over a commutative semiring, and the problem is to compute a (pre-) fixpoint of $f$, i.e. a point $x\in S^{n}$ where $x=f(x)$. As surveyed in much details in [10], this problem was studied in a very wide range of communities, such as in automata theory [14], program analysis [4, 18], graph algorithms [3, 17, 16] communities since the 1970s. (See [21, 8, 15, 7, 23] and references thereof).
The first approach is to keep iterating until a fixpoint is reached; in different contexts, this has different names: the naive algorithm, Jacobi iteration, Gauss-Seidel iteration, or Kleene iteration. The main advantage of this approach, which is studied in this paper, is that it assumes less about the underlying algebraic structure: we do not need both left and right distributive law, and do not need to assume a closure operator. The second approach is based on Gaussian elimination (Gauss-Jordan elimination), which, assuming oracle access to the solution $x^{*}$ of the 1D problem $x=1+ax$, can solve the algebraic path problem in $O(n^{3})$-time [21, 15]. The third approach is based on specifying the solutions based on the free semiring generated when viewing $\bm{A}$ as the adjacency matrix of a graph [22]. The underlying graph structure (such as planarity) may be exploited for very efficient algorithms [17, 16].
We also discuss the previous work on the fixed point solution to $x=f(x)$ in the non-linear case. See [20, 19, 13] for the study when $f$’s component functions are multivariate polynomials over the Kleene algebra. In the non-linear case, Kleene iteration may not converge in a finite number of steps. Thus, methods based on Galois connection or on widening/narrowing approaches [5] were proposed. These approaches are (discrete) lattice-theoretic. More recently, a completely different approach drawing inspiration from Newton’s method for solving a system of (real) equations was proposed [8, 6]. Newton’s method converges at least as fast as Kleene iteration, but each iteration could be more costly it need to compute a matrix inverse [6]. Recently, semi-naïve evaluation has been extended for a higher-order functional language called Datafun [2]. The book [7] contains many fundamental results on algebraic structures related to semirings and computational problems on such structures.
There are a daunting number of different statements, proofs and extensions of Parikh’s theorem in the literature. See for example [11]. As far as we know our extension of Parikh’s theorem does not appear in the literature, but given the volume of the work in this area, it is hard to be totally confident about the novelty of our extension.
Appendix A Semirings
In this section, we introduce the notation and terminology used throughout the paper.
Definition A.1 (Semiring).
A semiring [7] is a tuple $\bm{S}=(S,\oplus,\otimes,0,1)$ where
  • $\oplus$ and $\otimes$ are binary operators on $S$,
  • $(S,\oplus,0)$ is a commutative monoid, meaning $\oplus$ is commutative and associative, and 00 is the identity for $\oplus$,
  • $(S,\otimes,1)$ is a monoid, meaning $\otimes$ is associative, and 00 is the identity for $\oplus$,
  • 00 annilates every element $a\in S$, that is $a\otimes 0=0\otimes a=0$, and
  • $\otimes$ distributes over $\oplus$.
A commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ is a semiring where additionally $\otimes$ is commutative.
In the following fact about $p$-stable commutative semiring will be useful.
Proposition A.2.
Given a $p$-stable commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$, for any $u\in S$, we have $pu=(p+1)u$, where $pu$ here is shorthand for $\oplus_{i=1}^{p}u$.
Proof.
This follows directly from the $p$-stability of 1, and the fact that $11^{2}=$ 1. ∎